OpenStudy (anonymous):

4 times the lesser of the two consecutive even intergers is 12 less than twice the greater #'s . what are the intergers, and how did you get them?

OpenStudy (anonymous):

4[x+(x+2)] = 2[(x+4)+(x+6)] - 12

OpenStudy (anonymous):

4(2x+2) = 2(2x+10) - 12 8x + 8 = 4x + 20 - 12 8x + 8 = 4x + 8 4x + 8 = 8 4x = 0 x = 0 I'm not sure if I'm right because I'm half sleep.. I don't think so xD.. Anyone care to assist me?

OpenStudy (anonymous):

i know that the answer is -4,-2 but im not quite sure how.